OpenStudy (anonymous):

simplify the expression 121^1/2

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (anonymous):

\[\frac{ 121^1 }{ 2 }\] right?

Nnesha (nnesha):

convert 1/2 to square root exponent rule \[\huge\rm x^\frac{ m }{ n } = \sqrt[n]{x^m}\]

OpenStudy (anonymous):

ok thank you :D

Nnesha (nnesha):

so what did you get ?>

OpenStudy (anonymous):

um 121